Can you help me with this?

A= h (B+b);B
_
2

Answer by midwood_trail(310)

Can you help me with this?

A = h/2(B + b)...Solve for B
We can write h/2(B + b) as a single number over 2 this way:
A = [h(B+b)]/2
Apply the distributive property in the numerator.
A = (hB + hb)/2
We now have to remove the fraction on the right side.
We do this by multiplying both sides by 2.
2A = hB + hb
Now subtract hb from both sides of the equation.
Why are we doing all of this?
The goal is to ISOLATE B. This is what is meant by solve for B.
2A - hb = hB
Notice that we are solving for upper case B not lower case b.
Next and finally, divide both sides of the equation by h to find B.
(2A - hb)/h = B
We can reduce the left side a little more and get our final answer:
(2A/h) - b = B
Did you follow?